OpenStudy (anonymous):

NEED HELP FAST! I WILL GIVE MEDAL AND FAN!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

hartnn (hartnn):

what have you tried? :)

hartnn (hartnn):

could you find the slope of 3x - y = -3 ?

OpenStudy (pawanyadav):

Slope of line must be 3. y=3x+3 Comparwith y=mx+c

OpenStudy (pawanyadav):

Let eq be y=3x+c Find c by satisfying the equation with (4,-2) Then put value of c in above equation you will get the answer.

OpenStudy (pawanyadav):

Slope of line must be 3. y=3x+3 Comparwith y=mx+c

OpenStudy (anonymous):

So i have to change the equation to y=mx+b?

OpenStudy (pawanyadav):

Yes

OpenStudy (anonymous):

so it would be y=3x-3?

OpenStudy (pawanyadav):

Y=3x+3

OpenStudy (anonymous):

in the answer choices they're looking for a fraction in the equation

hartnn (hartnn):

you got how the slope is 3 ?

OpenStudy (anonymous):

yes i get that part

OpenStudy (anonymous):

becaus b= slope in y=mx+b

hartnn (hartnn):

yes, so your answer will be of the form y - 3x +c now we have the point 4,-2 plug in x= 4, y= -2 in that

hartnn (hartnn):

in y= 3x+c **

OpenStudy (pawanyadav):

No m is the slope .

OpenStudy (anonymous):

-2 = 3*4 + c

OpenStudy (pawanyadav):

b is the point where the line cuts the y-axis

OpenStudy (anonymous):

yes the y-intercept sorry i got confused for a second i realize that now thank you

hartnn (hartnn):

-2 = 3*4 + c yes thats right, now what is c ?

OpenStudy (anonymous):

3?

OpenStudy (pawanyadav):

Welcome, now solve other part carefully.

hartnn (hartnn):

how? -2 = 12 +c subtract 12 from both sides to get c

OpenStudy (anonymous):

-14 = c

OpenStudy (anonymous):

-2 =3*4 -14

hartnn (hartnn):

yes, so your equation as of now is y = 3x -14 right ? now what are your choices?

OpenStudy (anonymous):

no right now i have -2 = 3*4 - 14

hartnn (hartnn):

we had y = mx +c the slope we got , m = 3 so we were y= 3x +c. now we need to just find the "c" part. for which we used the given point. and got c = -14

hartnn (hartnn):

and if we plug c in y = 3x+c, we get y = 3x -14 :)

OpenStudy (anonymous):

yes! but the answer choiches all have fractions in them

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

that is correct thanks so much for your help!

hartnn (hartnn):

most welcome ^_^ hallp to help :)

hartnn (hartnn):

*happy